# Flatten Graphs as Sequences: Transformers are Scalable Graph Generators

### v2

@@ 2.3 Sampling Algorithm for SENT
Thm. 2.15 offers a simple strategy to sample a causal and Hamiltonian SENT: one needs to traverse the graph and choose the neighborhood set as all neighbors of the current node that have been visited. The traversing strategy could be achieved through a random path sampling or a depth-first search. In Algorithm 1, we provide a sampling strategy based on random path sampling with breaks.
Complexity analysis.
The length of a SENT, including the sizes of neighborhood sets (in other words, tokenized SENT defined in Section 2.4), is bounded by the number of nodes and edges, as each node and edge can be visited exactly once. Therefore, both the time and space complexity of sampling a SENT from graph $G$ are $\mathcal{O}(m)$ where $m$ is the number of edges.
@@ 3 Related Work
Autoregressive models generate graphs by sequentially adding nodes and edges. GraphRNN [66] pioneered this approach by framing graph generation as a sequence prediction task, demonstrating the capacity of recurrent neural networks (RNNs) [13] to capture complex structures. DeepGMG [38] introduced a probabilistic policy framework for conditional generation, while GRAN [39] and BiGG [14] enhanced efficiency and scalability by generating multiple nodes and edges in parallel.Recent research has focused on optimizing the generation order. Chen et al. [10] highlighted that the ordering of node and edge additions impacts graph quality, and GraphARM [36] applied reinforcement learning to dynamically refine this order. Goyal et al. [20] incorporated logical constraints to improve domain-specific generation, and Bacciu et al. [1] proposed Bayesian reasoning to better capture graph dependencies. BwR [16] and GEEL [29] investigated node ordering based on optimized bandwidth.These models, while efficient on synthetic datasets, do not explicitly represent graphs as token sequences, preventing direct application of LLM techniques. More significantly, their sequences are not guaranteed to be subgraph-induced (see Thm. 2.14). In contrast, our approach enables substructured-conditioned generation analogous to prompt-based generation in language models, establishing a more fundamental connection between graph and language modeling paradigms.
Other graph generative models.
Other graph generative models include variational, GAN-based, flow-based, and diffusion-based approaches. GraphVAEs [35, 53] employ variational autoencoders to learn latent representations, effectively generating small graphs but struggling with more complex structures. GAN-based models, such as NetGAN [4] and SPECTRE [41], generate graphs by modeling graph descriptors like random walks and spectral features. Flow-based methods such as [52] have shown the ability to generate small molecular graphs.Diffusion-based models iteratively refine noise into structured graphs through reverse diffusion steps. Continuous diffusion models [46, 31] adapt denoising diffusion probabilistic models for graph generation. To leverage graph sparsity and structure, discrete diffusion models [61, 36, 32, 64] have been developed. However, a key challenge for these models is the slow sampling process due to the long reverse diffusion chain. To mitigate this limitation, several efficient diffusion techniques have been proposed, including EDGE [11], HiGen [33], ESGG [2], and Pard [68].
Random walks for graph learning.
Random walks have been widely used in graph learning due to their strong expressive power. GCKN [8] and RWGNN [45] utilize path and walk kernels to learn graph representations. Several recent works [27, 62, 65] explicitly integrate random walk sequences with positional encodings, inspiring subsequent methods such as CRaWL [57], NeuralWalker [9] and RWNN [34], which further enhance graph representation learning via random walk sequence modeling. GraphGPT [69] leverages Eulerian paths to improve graph property prediction. Our work explores sequence representations of graphs for graph generation, introducing a novel perspective on combining random walks and language modeling for scalable graph generation.
@@ 4 Experiments
Implementation details.
We employ the LLaMA model with 12 layers and a hidden dimension of 768 as our sequence model backbone across all experiments, aligning with the architecture of GPT-2’s smallest variant [49]. Although prior works have used smaller models, we argue that our approach still demonstrates better scalability and faster training and inference speeds compared to diffusion models. For inference, we adopt the commonly used top-k sampling strategy [19]. Our implementation leverages the Hugging Face framework [28], providing users with a flexible interface to experiment with SOTA language models for graph generation.
Evaluation.
For fair comparison, we align our evaluation methodology with established practices from prior works [66, 41, 61]. Our evaluation compares generated samples against the test set using maximum mean discrepancy (MMD) [21], computed across multiple graph descriptors: node degree distributions (Deg.), clustering coefficients (Clus.), orbit count statistics (Orbit), and eigenvalue spectra (Spec.).As a reference, we also compute MMDs between the training and test sets and report the average ratio between generated and training MMDs (Ratio) following Bergmeister et al. [2].For synthetic datasets, we additionally assess model performance using the VUN metric, namely the proportion of generated graphs that are simultaneously valid, unique, and novel compared to the training graphs. Our efficiency analysis includes two measurements: inference speed, calculated as the per-graph generation time when producing 1024 graphs, and training speed, measured as the time required to achieve a VUN score of 75.0 for the Planar dataset and 60.0 for the SBM dataset. All efficiency measurements are performed on a single NVIDIA H100 GPU.For molecular generation datasets, we strictly follow the evaluation metrics used in DiGress [61] and use the evaluation tools from the official codebase [48, 6]. More details are provided in App. E.2.
@@ 4.1 Comparison to State-of-the-Art Methods
We evaluate the performance of AutoGraph compared to other SOTA graph generative models using the standard setting without pre-training.The comparison partners include GraphRNN [66], GRAN [39], SPECTRE [41], EDGE [11], GraphGen [20], BiGG [14], DiGress [61], GruM [32], GEEL [29], and ESGG [2].
Small synthetic graph generation.
Dataset Time DiGress Gran ESGG AutoGraph
[…]
We first evaluate our method on the small synthetic graph datasets introduced by Martinkus et al. [41], including the Planar and SBM datasets.As shown in Table 1, AutoGraph demonstrates competitive MMDs while ranking second-best and best in terms of VUN scores on the Planar and SBM datasets, respectively. Importantly, only GruM and AutoGraph exhibit strong structural validity (VUN $\geq 80.0$) on the SBM dataset. Previous state-of-the-art autoregressive models, particularly GEEL, completely fail on both datasets in terms of VUN scores, as they largely memorize some subset of the training data. It is worth noting that the relatively low MMD ratio of AutoGraph is expected, as we selected the best model based on the VUN score. More results with error bars are provided in App. F.1.Additionally, we assess the training and inference times of AutoGraph against representative models, including DiGress, GRAN, and ESGG. As presented in Table 2, AutoGraph is approximately 3 times faster during training and 100 times faster during inference compared to diffusion-based models. This substantial speedup over diffusion-based models is even more pronounced than that observed in other data modalities such as images [56].
@@
Large graph generation.
To understand the scalability of AutoGraph, we evaluate its performance on the Proteins and Point Clouds datasets used by Liao et al. [39]. The results, shown in Table 3, demonstrate that even when using a context window shorter than the longest sequence sampled from the dataset, AutoGraph achieves MMD ratios comparable to those observed on the Planar and SBM datasets. Furthermore, AutoGraph outperforms all existing methods in terms of MMD ratio, achieving a twofold or more improvement over the previous best model, ESGG. More significantly, while ESGG was specifically designed for generating unattributed graphs, AutoGraph demonstrates versatility by being applicable to both unattributed and attributed graphs. Finally, previous state-of-the-art autoregressive models, such as GEEL, again fail to achieve competitive performance on these datasets.
Molecular graph generation.
We demonstrate the applicability of our method to generating real-world attributed graphs, such as molecular structures. We evaluate AutoGraph on the same datasets used by DiGress [61], including QM9 (all atoms) [63], MOSES [48], and GuacaMol [6]. Following the data splits and experimental setup from DiGress, we benchmark AutoGraph against a variety of SOTA models, including DiGress, VAE on SMILES [48], JT-VAE [30], GraphINVENT [43], NAGVAE [37], LSTM and MCTS [6]. On the QM9 dataset (Table 12), AutoGraph outperforms DiGress across all metrics except uniqueness, showing its superiority for attributed graphs.For the more challenging MOSES and GuacaMol datasets, AutoGraph also demonstrates superior performance, achieving higher validity and improved distributional alignment as measured by metrics like FCD, as shown in Table 4. Notably, to our best knowledge, AutoGraph is the first autoregressive model for graphs to surpass diffusion-based approaches on these datasets. It is worth mentioning that all metrics were computed using SMILES representations rather than molecular graphs. Due to the non-reversible nature of converting SMILES to graphs and back, where approximately 20% of molecules cannot be mapped back to their original SMILES [61], some discrepancies are introduced when calculating these metrics. Despite these challenges, AutoGraph achieves validity and FCD scores comparable to SMILES-based methods.Furthermore, AutoGraph demonstrates remarkable efficiency, with training times of less than one day on both datasets, compared to up to one week for DiGress [61]. This substantial reduction underscores AutoGraph’s practical advantages in large-scale molecular graph generation tasks.
@@ 4.4 Ablation Experiments
In this study, we aim to understand the effectiveness of the key components in AutoGraph.
Comparison of sequence model architectures.
AutoGraph provides a novel framework for evaluating the capability of current LLM architectures in graph generation and, more broadly, in structural reasoning tasks. In Table 7, we compare several state-of-the-art architectures on the Planar dataset, including GPT-2 [49], Mamba [22], and LLaMA [58]. While all models achieve comparable MMD ratios, transformer-based architectures, particularly LLaMA, demonstrate significantly better performance in terms of VUN scores compared to state-space models. These findings highlight the potential of AutoGraph to serve as a valuable benchmark for assessing sequence/language models’ capabilities in graph generation tasks.
@@
Effect of top-k sampling.
A key advantage of AutoGraph over diffusion-based approaches is the flexibility to apply top-k sampling [19] during inference, which can improve generation quality. As shown on the left of Figure 2, a smaller $k$ improves the VUN score on the Planar dataset, whereas it is not beneficial on the SBM dataset. In contrast, increasing $k$ generally improves MMD ratios on both datasets. These results suggest that top-k sampling can be optimized based on dataset characteristics. In our experiments, we select the best $k$ that maximizes the VUN score for small synthetic datasets and minimizes the validation MMD ratios for other datasets. This flexibility allows practitioners to select $k$ based on specific performance criteria they aim to prioritize.
Comparison of SET and SENT.
As discussed in Section 2, SENT is preferred over SET for graph generation, as incorporating neighborhood information is essential to ensure structural coherence. To empirically validate this, we compare the performance of SENT and SET on the Planar dataset and present the training curves on the right of Figure 2. Consistent with our theoretical analysis, SET fails to produce high-validity graphs, resulting in a VUN score close to zero, whereas SENT successfully generates valid planar graphs.
@@ 5 Conclusion
We proposed AutoGraph, a scalable and efficient autoregressive model for attributed graph generation that handles large graphs while maintaining high quality. Our approach enables substructure-conditioned generation without additional fine-tuning and demonstrates promising transfer capabilities. Crucially, AutoGraph establishes the first fundamental connection between graph and language modeling–where graphs are losslessly represented as token sequences, and prefixes in these sequences serve as meaningful patterns in both paradigms–representing a significant step toward applying language modeling techniques to graph generation and broader graph learning challenges.
Limitations.
While AutoGraph demonstrates strong scalability on current graph generation benchmarks, we acknowledge that the datasets used in our study remain relatively small-scale compared to those used in pre-training LLMs. To push the boundaries of more powerful graph generative models or eventually foundation models for graphs, we draw the community’s attention to building more comprehensive graph generation benchmarks and well-curated pre-training datasets.
@@ Acknowledgements
The authors thank Dr. Till Hendrik Schulz, Philip Hartout, and Błażej Banaszewski for their insightful discussions and valuable feedback on the manuscript.
Appendix
Appendix A Background on Sequence Model Architectures
GPT-2.
LLaMA.
Mamba.
Appendix B Broader Impacts
Appendix C Additional Details about AutoGraph
C.1 Background on Graph Theory
Definition C.1 (Graph isomorphism).
Definition C.2 (Induced subgraph).
C.2 Remarks on Tokenized SENTs
[…]
C.3 Remarks on Model Inference
[…]
Appendix D Proofs
Appendix E Experimental Details
E.1 Datasets
Small synthetic graphs: Planar and SBM.
Large graphs: Proteins and Point Clouds.
QM9.
MOSES and GuacaMol.
PubChem-10M.
NetworkX.
Table 8: Dataset statistics
Train Val Test
Unattributed Graphs
Planar 128 32 40 64 64 181 178
SBM 128 32 40 187 104 1129 500
Proteins 587 147 184 500 258 1575 646
Point Clouds 26 7 8 5037 1332 10886 2971
Attributed Graphs
QM9 97734 20042 13055 29 18 28 19
MOSES 1584663 176225 176074 27 22 31 23
GuacaMol 1118633 69926 209654 88 28 88 30
Pre-training unattributed graphs
NetworkX 24957 2516 5999 459 5999 751
E.2 Evaluation Metrics
Planar and SBM.
QM9.
MOSES and GuacaMol.
  • Validity: Proportion of molecules passing basic valency checks.
E.3 Computing Details
E.4 Hyperparameters
Table 10: Inference hyperparameters for each dataset.
w/o pre-training w/ pre-training
Dataset Top-$k$ Temperature $\tau$ Top-$k$ Temperature $\tau$
Planar 10 1.0 30 0.9
SBM 60 1.0 150 1.0
Proteins 40 1.0 30 1.05
Point Clouds 60 1.0 20 0.9
NetworkX 120 1.0
QM9 5 1.0
MOSES 5 1.0
GuacaMol 5 1.0
Appendix F Additional Results
F.1 Additional Results on Synthetic Datasets
Table 11: Performances on the Planar and SBM datasets with error bars
(a) Planar
Model Deg. Clus. Orbit Spec. Ratio VUN
(b) SBM
Model Deg. Clus. Orbit Spec. Ratio VUN
F.2 Additional Results on molecular generation datasets
Table 12: Benchmarking AutoGraph on the QM9 dataset
QM9 with hydrogen atoms
[…]
DiGress 95.4 97.6 33.4 98.1 79.8
AutoGraph 97.7 96.7 45.5 98.6 87.3
Table 13: Benchmarking AutoGraph on the MOSES dataset
MOSES
[…]
VAE SMILES 97.7 99.8 69.5 99.7 0.57 0.58 5.9
JT-VAE Fragments 100 100 99.9 97.8 1.00 0.53 10
GraphINVENT Graph 96.4 99.8 95.0 1.22 0.54 12.7
DiGress Graph 85.7 100 95.0 97.1 1.19 0.52 14.8
AutoGraph Graph 87.4 100 85.9 98.6 0.91 0.55 10.2
F.3 Transfer Performance of AutoGraph
F.4 Substructure Conditioned Generation
F.5 Additional Ablation Experiments
Impact of model size.
Table 14: Comparison of model size on the Planar dataset
Model # Params Configuration Deg. Clus. Orbit Spec. Ratio VUN
LLaMA-s 113M 12 layers, 768 dims 0.0005 0.0651 0.0005 0.0056 1.6 90.0
LLaMA-m 402M 24 layers, 1024 dims 0.0001 0.0340 0.0013 0.0064 1.4 82.5
F.6 Visualization of Graphs Generated by AutoGraph
F.6.1 Results without Pre-training
F.6.2 Results with pre-training

### v3

@@ 2.3 Sampling Algorithm for SENT
Thm. 2.15 offers a simple strategy to sample a causal and Hamiltonian SENT: one needs to traverse the graph and choose the neighborhood set as all neighbors of the current node that have been visited. The traversing strategy could be achieved through a random path sampling or a depth-first search. In Algorithm 1, we provide a sampling strategy based on random path sampling with breaks.Complexity analysis.The length of a SENT, including the sizes of neighborhood sets (in other words, tokenized SENT defined in Section 2.4), is bounded by the number of nodes and edges, as each node and edge can be visited exactly once. Therefore, both the time and space complexity of sampling a SENT from graph $G$ are $\mathcal{O}(m)$ where $m$ is the number of edges.
@@ 3 Related Work
Autoregressive models for graph generation.Autoregressive models generate graphs by sequentially adding nodes and edges. GraphRNN [71] pioneered this approach by framing graph generation as a sequence prediction task, demonstrating the capacity of recurrent neural networks (RNNs) [15] to capture complex structures. DeepGMG [42] introduced a probabilistic policy framework for conditional generation, while GRAN [43] and BiGG [16] enhanced efficiency and scalability by generating multiple nodes and edges in parallel. ANFM [40] leverages filtration to improve efficiency.Recent research has focused on optimizing the generation order. Chen et al. [12] highlighted that the ordering of node and edge additions impacts graph quality, and GraphARM [39] applied reinforcement learning to dynamically refine this order. Goyal et al. [23] incorporated logical constraints to improve domain-specific generation, and Bacciu et al. [1] proposed Bayesian reasoning to better capture graph dependencies. BwR [18] and GEEL [32] investigated node ordering based on optimized bandwidth.These models, while efficient on synthetic datasets, do not explicitly represent graphs as token sequences, preventing direct application of LLM techniques. More significantly, their sequences are not guaranteed to be subgraph-induced (see Thm. 2.14). In contrast, our approach enables substructured-conditioned generation analogous to prompt-based generation in language models, establishing a more fundamental connection between graph and language modeling paradigms.Other graph generative models.Other graph generative models include variational, GAN-based, flow-based, and diffusion-based approaches. GraphVAEs [38, 58] employ variational autoencoders to learn latent representations, effectively generating small graphs but struggling with more complex structures. GAN-based models, such as NetGAN [4] and SPECTRE [45], generate graphs by modeling graph descriptors like random walks and spectral features. Flow-based methods such as [57] have shown the ability to generate small molecular graphs.Diffusion-based models iteratively refine noise into structured graphs through reverse diffusion steps. Continuous diffusion models [51, 34] adapt denoising diffusion probabilistic models for graph generation. To leverage graph sparsity and structure, discrete diffusion models [66, 39, 35, 69] have been developed. However, a key challenge for these models is the slow sampling process due to the long reverse diffusion chain. To mitigate this limitation, several efficient diffusion techniques have been proposed, including EDGE [13], HiGen [36], ESGG [2], and Pard [73].Random walks for graph learning.Random walks have been widely used in graph learning due to their strong expressive power. GCKN [8] and RWGNN [50] utilize path and walk kernels to learn graph representations. Several recent works [30, 67, 70] explicitly integrate random walk sequences with positional encodings, inspiring subsequent methods such as CRaWL [62], NeuralWalker [10] and RWNN [37]. GraphGPT [74] leverages Eulerian paths to improve graph property prediction.Some graph transformers [48, 9] also leverage features based on random walks. Moreover, graph-to-sequence representations have been used to assist LLMs in understanding graphs [22, 11].Our work explores sequence representations of graphs for graph generation, introducing a novel perspective on combining random walks and language modeling for scalable graph generation.
@@ 4 Experiments
Implementation details.We employ the LLaMA model with 12 layers and a hidden dimension of 768 as our sequence model backbone across all experiments, aligning with the architecture of GPT-2’s smallest variant [54]. Although prior works have used smaller models, we argue that our approach still demonstrates better scalability and faster training and inference speeds compared to diffusion models. For inference, we adopt the commonly used top-k sampling strategy [21]. Our implementation leverages the Hugging Face framework [31], providing users with a flexible interface to experiment with SOTA language models for graph generation.Evaluation.For fair comparison, we align our evaluation methodology with established practices from prior works [71, 45, 66]. Our evaluation compares generated samples against the test set using maximum mean discrepancy (MMD) [24], computed across multiple graph descriptors: node degree distributions (Deg.), clustering coefficients (Clus.), orbit count statistics (Orbit), and eigenvalue spectra (Spec.).As a reference, we also compute MMDs between the training and test sets and report the average ratio between generated and training MMDs (Ratio) following Bergmeister et al. [2].For synthetic datasets, we additionally assess model performance using the VUN metric, namely the proportion of generated graphs that are simultaneously valid, unique, and novel compared to the training graphs. Our efficiency analysis includes two measurements: inference speed, calculated as the per-graph generation time when producing 1024 graphs, and training speed, measured as the time required to achieve a VUN score of 75.0 for the Planar dataset and 60.0 for the SBM dataset. All efficiency measurements are performed on a single NVIDIA H100 GPU.For molecular generation datasets, we strictly follow the evaluation metrics used in DiGress [66] and use the evaluation tools from the official codebase [53, 6]. More details are provided in App. E.2.
@@ 4.1 Comparison to State-of-the-Art Methods
We evaluate the performance of AutoGraph compared to other SOTA graph generative models using the standard setting without pre-training.The comparison partners include GraphRNN [71], GRAN [43], SPECTRE [45], EDGE [13], GraphGen [23], BiGG [16], DiGress [66], GruM [35], GEEL [32], and ESGG [2].Small synthetic graph generation.We first evaluate our method on the small synthetic graph datasets introduced by Martinkus et al. [45], including the Planar and SBM datasets.As shown in Table 1, AutoGraph demonstrates competitive MMDs while ranking second-best and best in terms of VUN scores on the Planar and SBM datasets, respectively. Importantly, only GruM and AutoGraph exhibit strong structural validity (VUN $\geq 80.0$) on the SBM dataset. Previous state-of-the-art autoregressive models, particularly GEEL, completely fail on both datasets in terms of VUN scores, as they largely memorize some subset of the training data. It is worth noting that the relatively low MMD ratio of AutoGraph is expected, as we selected the best model based on the VUN score. More results with error bars are provided in App. F.1.Additionally, we assess the training and inference times of AutoGraph against representative models, including DiGress, GRAN, and ESGG. As presented in Table 3, AutoGraph is approximately 3 times faster during training and 100 times faster during inference compared to diffusion-based models. This substantial speedup over diffusion-based models is even more pronounced than that observed in other data modalities such as images [61].
@@
Dataset Time DiGress Gran ESGG AutoGraph
[…]
Large graph generation.To understand the scalability of AutoGraph, we evaluate its performance on the Proteins and Point Clouds datasets used by Liao et al. [43]. The results, shown in Table 2, demonstrate that even when using a context window shorter than the longest sequence during training, AutoGraph achieves MMD ratios comparable to those observed on the Planar and SBM datasets. Furthermore, AutoGraph outperforms all existing methods in terms of MMD ratio, achieving a twofold or more improvement over the previous best model, ESGG. More significantly, while ESGG was specifically designed for generating unattributed graphs, AutoGraph demonstrates versatility by being applicable to both unattributed and attributed graphs. Finally, previous state-of-the-art autoregressive models, such as GEEL, again fail to achieve competitive performance on these datasets.Molecular graph generation.We demonstrate the applicability of our method to generating real-world attributed graphs, such as molecular structures. We evaluate AutoGraph on the same datasets used by DiGress [66], including QM9 (all atoms) [68], MOSES [53], and GuacaMol [6]. Following the data splits and experimental setup from DiGress, we benchmark AutoGraph against a variety of SOTA models, including DiGress, VAE on SMILES [53], JT-VAE [33], GraphINVENT [47], NAGVAE [41], LSTM and MCTS [6]. On the QM9 dataset (Table 12), AutoGraph outperforms DiGress across all metrics except uniqueness, showing its superiority for attributed graphs.For the more challenging MOSES and GuacaMol datasets, AutoGraph also demonstrates superior performance, achieving higher validity and improved distributional alignment as measured by metrics like FCD, as shown in Table 4. Notably, to our best knowledge, AutoGraph is the first autoregressive model for graphs to surpass diffusion-based approaches on these datasets. It is worth mentioning that all metrics were computed using SMILES representations rather than molecular graphs. Due to the non-reversible nature of converting SMILES to graphs and back, where approximately 20% of molecules cannot be mapped back to their original SMILES [66], some discrepancies are introduced when calculating these metrics. Despite these challenges, AutoGraph achieves validity and FCD scores comparable to SMILES-based methods.Furthermore, AutoGraph demonstrates remarkable efficiency, with training times of less than one day on both datasets, compared to up to one week for DiGress [66]. This substantial reduction underscores AutoGraph’s practical advantages in large-scale molecular graph generation tasks.
@@ 4.4 Ablation Experiments
In this study, we aim to understand the effectiveness of the key components in AutoGraph.Comparison of sequence model architectures.AutoGraph provides a novel framework for evaluating the capability of current LLM architectures in graph generation and, more broadly, in structural reasoning tasks. In Table 7, we compare several state-of-the-art architectures on the Planar dataset, including GPT-2 [54], Mamba [25], and LLaMA [63]. While all models achieve comparable MMD ratios, transformer-based architectures, particularly LLaMA, demonstrate significantly better performance in terms of VUN scores compared to state-space models. These findings highlight the potential of AutoGraph to serve as a valuable benchmark for assessing sequence/language models’ capabilities in graph generation tasks.
@@
Effect of top-k sampling.A key advantage of AutoGraph over diffusion-based approaches is the flexibility to apply top-k sampling [21] during inference, which can improve generation quality. As shown on the left of Figure 2, a smaller $k$ improves the VUN score on the Planar dataset, whereas it is not beneficial on the SBM dataset. In contrast, increasing $k$ generally improves MMD ratios on both datasets. These results suggest that top-k sampling can be optimized based on dataset characteristics. In our experiments, we select the best $k$ that maximizes the VUN score for small synthetic datasets and minimizes the validation MMD ratios for other datasets. This flexibility allows practitioners to select $k$ based on specific performance criteria they aim to prioritize.Comparison of SET and SENT.As discussed in Section 2, SENT is preferred over SET for graph generation, as incorporating neighborhood information is essential to ensure structural coherence. To empirically validate this, we compare the performance of SENT and SET on the Planar dataset and present the training curves on the right of Figure 2. Consistent with our theoretical analysis, SET fails to produce high-validity graphs, resulting in a VUN score close to zero, whereas SENT successfully generates valid planar graphs.
@@ 5 Conclusion
We proposed AutoGraph, a scalable and efficient autoregressive model for attributed graph generation that handles large graphs while maintaining high quality. Our approach enables substructure-conditioned generation without additional fine-tuning and demonstrates promising transfer capabilities. Crucially, AutoGraph establishes the first fundamental connection between graph and language modeling–where graphs are losslessly represented as token sequences, and prefixes in these sequences serve as meaningful patterns in both paradigms–representing a significant step toward applying language modeling techniques to graph generation and broader graph learning challenges.Limitations.While AutoGraph demonstrates strong scalability on current graph generation benchmarks, we acknowledge that the datasets used in our study remain relatively small-scale compared to those used in pre-training LLMs. To push the boundaries of more powerful graph generative models or eventually foundation models for graphs, we draw the community’s attention to building more comprehensive graph generation benchmarks and well-curated pre-training datasets.
@@ Acknowledgements
The authors thank Dr. Till Hendrik Schulz, Philip Hartout, and Błażej Banaszewski for their insightful discussions and valuable feedback on the manuscript. The authors thank Dr. Till Hendrik Schulz for presenting the poster at NeurIPS, as the first author was unable to attend due to a visa issue.